A pair of narrow slits that are 1.8 mm apart is illuminated by a monochromatic coherent light source. A fringe pattern is observed on a screen 4.8 m from the slits. If there are 5.0 bright fringes/cm on the screen, what is the wavelength of the monochromatic light?

A) 550 nm       B) 600 nm       C) 650 nm       D) 700 nm       E) 750 nm
Answer #1

apply Y = mLR/d

d = 1.8 mm

so wavelgnth L = Yd/mR

L = (1*10^-2 * 1.8 mm/(5 * 4.8)

L = 750 nm